OpenStudy (anonymous):

16x + -3(4x + 7) = 6x + -1(2x + 21) for x

OpenStudy (maheshmeghwal9):

wt u have done up-to now? Show me some steps i will help u further:)

OpenStudy (anonymous):

well i have this so far- 16x-12x-21=6x-2x+21 4x-21=4x+21

OpenStudy (maheshmeghwal9):

well u have made a mistake:/

OpenStudy (maheshmeghwal9):

-1(2x+21)= -2x-21.

OpenStudy (anonymous):

where ?

OpenStudy (anonymous):

i dont understand

OpenStudy (maheshmeghwal9):

well i m copying & paste ur reply with correction when u gt ur mistake tell me

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

well how is 21 negative on the rhs

OpenStudy (maheshmeghwal9):

oops sorry wait a minute: - this is correct 16x-12x-21=6x-2x-21 4x-21=4x-21

OpenStudy (maheshmeghwal9):

so how negative 21.

OpenStudy (maheshmeghwal9):